by using row reduction solve the following simultaneous equation.
2x-y=5
3x+4y=13

Answer by checkley77(12844)

2x-y=5 multiply by 4 & add.
3x+4y=13
8x-4y=20
---------------
11x=33
x=3/11
x=3 ans.
2*3-y=5
6-y=5
-y=5-6
-y=-1
y=1 ans.
Proof:
3*3+4*1=13
9+4=13
13=13